# Infinite-Dimensional Stochastic Differential Equations and Diffusion Dynamics of Coulomb Random Point Fields

### v3

@@ 5.1. Construction of the logarithmic derivative $\mathsf{d}_{R,\mathsf{y}}$ of $\mu_{R,\mathsf{y}}$
$\lim_{N\to\infty}\pi_{\overline{S}_{R}^{c}}(\mathsf{y}^{N})=\pi_{\overline{S}_{R}^{c}}(\mathsf{y})\quad\text{ for all $\mathsf{y}\in\mathsf{S}$}.$
From ((A1)) and ((A2)), the labeled density $\mathfrak{m}^{N}$ is a continuous function on $O_{N}^{N}$. Hence, $\mu_{R,\mathsf{y}^{N}}^{N}$ can be defined for all $\mathsf{y}\in\mathsf{S}$ such that $\mathsf{y}(O_{N})=N$ and $\mathsf{y}(\mathbb{R}^{d}\backslash O_{N})=0$.
Using (2.18), we obtain the following for $\mu^{N}$-a.s. $\mathsf{y}$ and $({x},\mathsf{s})\in\overline{S}_{R}\times\overline{\mathsf{S}}_{R}$
$\mathsf{d}_{R,\mathsf{y}}^{N}({x},\mathsf{s})=$ $\mathsf{d}^{N}(x,\mathsf{s}+\pi_{\overline{S}_{R}^{c}}(\mathsf{y}))$
$=$ $-\beta\Big{\{}\nabla\Phi^{N}({x})+\sum_{s^{i}\in\overline{S}_{R}}\nabla_{{x}}\Psi^{N}({x},s^{i})+\sum_{i=1,\,{y}^{i}\not\in\overline{S}_{R}}^{N}\nabla_{{x}}\Psi^{N}({x},{y}^{i})\Big{\}},$ (5.4)
where we label $\mathsf{y}=\sum_{i}\delta_{{y}^{i}}$ such that ${y}^{i}=\mathfrak{l}^{i}(\mathsf{y})$. Then, from (2.43), we have $\lvert{y}^{i}\rvert\leq\lvert{y}^{i+1}\rvert$ for all ii. For $\mathsf{y}$ satisfying $\mathsf{y}(\overline{S}_{R})\geq 1$, (5.3) and (5.4) imply that, for $({x},\mathsf{s})\in\overline{S}_{R}\times\overline{\mathsf{S}}_{R}$,
$\mathsf{d}_{R,\mathsf{y}^{N}}^{N}({x},\mathsf{s})$ $=-\beta\Big{\{}\nabla\Phi^{N}({x})+\sum_{s^{i}\in\overline{S}_{R}}\nabla_{{x}}\Psi^{N}({x},s^{i})+\sum_{i=1,\,{y}^{i}\not\in\overline{S}_{R}}^{N}\nabla_{{x}}\Psi^{N}({x},{y}^{i})\Big{\}}.$ (5.5)
Let $a_{\mathsf{r}}=\{a_{\mathsf{r}}(R)\}_{R\in\mathbb{N}}$ be an increasing sequence of increasing sequences of natural numbers. Let $\mathsf{K}[a_{\mathsf{r}}]$ be the set defined by (2.44). Then, $\{\mathsf{K}[a_{\mathsf{r}}]\}_{\mathsf{r}\in\mathbb{N}}$ is an increasing sequence of compact sets. From (2.45), $\mu^{N_{\mathfrak{n}}}(\cup_{\mathsf{r}\in\mathbb{N}}\mathsf{K}[a_{\mathsf{r}}])=1$. Let
$\mu_{R,\mathsf{y}^{N_{\mathfrak{n}}},\mathsf{r}}^{N_{\mathfrak{n}}}=\mu_{R,\mathsf{y}^{N_{\mathfrak{n}}}}^{N_{\mathfrak{n}}}(\cdot\cap\mathsf{K}[a_{\mathsf{r}}]).$ (5.6)
Let $\mu_{R,\mathsf{y}^{N_{\mathfrak{n}}},\mathsf{r}}^{N_{\mathfrak{n}},[1]}$ be the one-reduced Campbell measure of $\mu_{R,\mathsf{y}^{N_{\mathfrak{n}}},\mathsf{r}}^{N_{\mathfrak{n}}}$. Let $\mathsf{f}_{Q}$ be as in (4.5) and $Q<R\in\mathbb{N}$. For $\mathsf{y}\in\cup_{\mathsf{r}=1}^{\infty}\mathsf{K}[a_{\mathsf{r}}]$, let $\mathsf{r}\in\mathbb{N}$ be such that $\mathsf{y}\in\mathsf{K}[a_{\mathsf{r}}]\backslash\mathsf{K}[a_{\mathsf{r}-1}]$. Then, for $\mathsf{y}\in\cup_{\mathsf{r}=1}^{\infty}\mathsf{K}[a_{\mathsf{r}}]$ and $Q<R,\mathfrak{n}\in\mathbb{N}$, we define the constants $c_{5.1}$, $c_{5.2}$, and $c_{5.3}$ by
@@
From (2.44) and (5.6), we have, for all $Q<R,\mathfrak{n}\in\mathbb{N}$, and $\mathsf{y}\in\mathsf{S}$,
$\int_{\mathsf{S}}\mathsf{s}(S_{Q+1})\mu_{R,\mathsf{y}^{N_{\mathfrak{n}}},\mathsf{r}}^{N_{\mathfrak{n}}}(d\mathsf{s})\leq a_{\mathsf{r}}(Q+1).$ (5.8)
Hence, from (5.7) and (5.8), we obtain, for $\mathsf{y}\in\mathsf{K}[a_{\mathsf{r}}]\backslash\mathsf{K}[a_{\mathsf{r}-1}]$,
@@
Hence, from ((A1)), (4.5), (5.7), and (5.9), we obtain Lemma 5.1. ∎
The following result is a local, quench version of Lemma 4.3 and Lemma 4.5.
Lemma 5.2.
Let $\mathsf{f}_{Q}$ be as in (4.5). Let $R\geq Q+1$. Then, for $\mathsf{y}\in\cup_{\mathsf{r}\in\mathbb{N}}\mathsf{K}[a_{\mathsf{r}}]$,
[…]
Proof.
We prove the strong convergence of the density function of $\mu_{R,\mathsf{y}^{N_{\mathfrak{n}}}}^{N_{\mathfrak{n}}}$, where $\mu_{R,\mathsf{y}^{N_{\mathfrak{n}}}}^{N_{\mathfrak{n}}}$ is defined by (5.1) for $N_{\mathfrak{n}}$ and $\mathsf{y}^{N_{\mathfrak{n}}}$. Let $\mu_{R,\mathsf{y}^{N_{\mathfrak{n}}}}^{N_{\mathfrak{n}},[1]}$ be the one-reduced Campbell measure of $\mu_{R,\mathsf{y}^{N_{\mathfrak{n}}}}^{N_{\mathfrak{n}}}$. Let $\mathfrak{m}_{R,\mathsf{y}^{N_{\mathfrak{n}}}}^{N_{\mathfrak{n}},[1]}$ be the Radon–Nikodym density of $\mu_{R,\mathsf{y}^{N_{\mathfrak{n}}}}^{N_{\mathfrak{n}},[1]}$ with respect to $\Lambda_{R}^{[1]}$, where $\Lambda_{R}^{[1]}$ was defined as in Section 4. Such densities $\mathfrak{m}_{R,\mathsf{y}^{N_{\mathfrak{n}}}}^{N_{\mathfrak{n}},[1]}$ exist from (2.18).
Let $\overline{\mathsf{S}}_{R,k}^{[1]}$ be as in (4.16). We set
$\mathfrak{m}_{R,\mathsf{y}^{N_{\mathfrak{n}}},k,l}^{N_{\mathfrak{n}},[1]}=\big{(}\mathfrak{m}_{R,\mathsf{y}^{N_{\mathfrak{n}}}}^{N_{\mathfrak{n}},[1]}1_{\overline{\mathsf{S}}_{R,k}^{[1]}}\big{)}\wedge l.$ (5.10)
By definition, $\mu_{R,\mathsf{y}}=\mu(\cdot|\pi_{\overline{S}_{R}^{c}}(\mathsf{s})=\pi_{\overline{S}_{R}^{c}}(\mathsf{y}))$ is a random point field on $\mathbb{R}^{d}$ concentrated at $\pi_{\overline{S}_{R}^{c}}(\mathsf{y})$ on $\overline{S}_{R}^{c}$. We regard $\mu_{R,\mathsf{y}}$ as a random point field on $\overline{S}_{R}$ as well as that on $\mathbb{R}^{d}$. Let $\mu_{R,\mathsf{y}}^{[1]}$ be the one-reduced Campbell measure of $\mu_{R,\mathsf{y}}$.
Lemma 5.3.
Let $R\geq Q+1$, $k\in\{0\}\cup\mathbb{N}$, and $l\in\mathbb{N}$. For $\mu$-a.s. $\mathsf{y}$, the following hold.
i $\{\mathsf{f}_{Q}\mathfrak{m}_{R,\mathsf{y}^{N_{\mathfrak{n}}},k,l}^{N_{\mathfrak{n}},[1]}\}_{\mathfrak{n}\in\mathbb{N}}$ is relatively compact in $L^{2}(\Lambda_{R}^{[1]})$.
$i$ $\mu_{R,\mathsf{y}}^{[1]}$ has a Radon-Nikodym density $\mathfrak{m}_{R,\mathsf{y}}^{[1]}$ with respect to $\Lambda_{R}^{[1]}$ satisfying
$\lim_{\mathfrak{n}\to\infty}\mathsf{f}_{Q}\mathfrak{m}_{R,\mathsf{y}^{N_{\mathfrak{n}}},k,l}^{N_{\mathfrak{n}},[1]}=\mathsf{f}_{Q}\big{(}\mathfrak{m}_{R,\mathsf{y}}^{[1]}1_{\overline{\mathsf{S}}_{R,k}^{[1]}})\wedge l\text{ in }L^{2}(\Lambda_{R}^{[1]}).$
Proof.
Lemma 5.3 follows from Lemma 5.2 using similar arguments as for the proof of Lemma 4.6 and Theorem 4.1. Hence, the proof is omitted here. ∎
Let $\overline{\mathsf{S}}_{R,k}^{[1]}$, $\mathfrak{m}_{R,\mathsf{y}}^{N,[1]}$, and $\mathfrak{m}_{R,\mathsf{y},k,l}^{N,[1]}$ be as in (5.10). Let
$\mathsf{M}_{R,\mathsf{y},k,l}^{N,[1]}$ $=\{({x},\mathsf{s})\in\overline{\mathsf{S}}_{R,k}^{[1]};0<\mathfrak{m}_{R,\mathsf{y}}^{N,[1]}({x},\mathsf{s})<l\},$
$\mathsf{d}_{R,\mathsf{y},k,l}^{N}({x},\mathsf{s})$ $=1_{\mathsf{M}_{R,\mathsf{y},k,l}^{N,[1]}}({x},\mathsf{s})\nabla_{{x}}\log\mathfrak{m}_{R,\mathsf{y},k,l}^{N,[1]}({x},\mathsf{s}).$
Let $\mathfrak{m}_{R,\mathsf{y},k,l}^{[1]}=\big{(}\mathfrak{m}_{R,\mathsf{y}}^{[1]}1_{\overline{\mathsf{S}}_{R,k}^{[1]}})\wedge l$. We define $\mathsf{M}_{R,\mathsf{y},k,l}^{[1]}$ and ${\mathsf{d}}_{R,\mathsf{y},k,l}$ by
$\mathsf{M}_{R,\mathsf{y},k,l}^{[1]}$ $=\{({x},\mathsf{s})\in\overline{\mathsf{S}}_{R,k}^{[1]};0<\mathfrak{m}_{R,\mathsf{y}}^{[1]}({x},\mathsf{s})<l\},$
${\mathsf{d}}_{R,\mathsf{y},k,l}({x},\mathsf{s})$ $=1_{\mathsf{M}_{R,\mathsf{y},k,l}^{[1]}}({x},\mathsf{s})\nabla_{{x}}\log\mathfrak{m}_{R,\mathsf{y},k,l}^{[1]}({x},\mathsf{s}).$ (5.11)
Lemma 5.4.
For $R\geq Q+1$, $k\in\{0\}\cup\mathbb{N}$, $l\in\mathbb{N}$, and $\mu$-a.s. $\mathsf{y}$,
Proof.
Lemma 5.4 follows from Lemma 5.2 and Lemma 5.3 using similar arguments as for the proof of Lemma 4.7. Hence, the proof is omitted here. ∎
Let $\mathsf{M}_{R,\mathsf{y}}^{[1]}=\{({x},\mathsf{s})|0<\mathfrak{m}_{R,\mathsf{y}}^{[1]}({x},\mathsf{s})<\infty\}$ and $\mathsf{d}_{R,\mathsf{y}}({x},\mathsf{s})=1_{\mathsf{M}_{R,\mathsf{y}}^{[1]}}({x},\mathsf{s})\nabla_{{x}}\log\mathfrak{m}_{R,\mathsf{y}}^{[1]}({x},\mathsf{s})$. From (5.11), we derive the following:
$\mathsf{d}_{R,\mathsf{y}}({x},\mathsf{s}):=\lim_{l\to\infty}\sum_{k=0}^{\infty}\mathsf{d}_{R,\mathsf{y},k,l}({x},\mathsf{s}),\quad\sum_{k=0}^{\infty}\lvert\mathsf{d}_{R,\mathsf{y},k,l}({x},\mathsf{s})\rvert\leq\lvert\mathsf{d}_{R,\mathsf{y}}({x},\mathsf{s})\rvert$. (5.12)
Lemma 5.5.
For all $R\geq Q+1$ and $\mu$-a.s. $\mathsf{y}$,
[…]
$\lim_{\mathfrak{n}\to\infty}\mathsf{d}_{R,\mathsf{y}^{N_{\mathfrak{n}}}}^{N_{\mathfrak{n}}}(\mathsf{f}_{Q}\mathfrak{m}_{R,\mathsf{y}^{N_{\mathfrak{n}}}}^{N_{\mathfrak{n}},[1]})^{1/2}=\mathsf{d}_{R,\mathsf{y}}(\mathsf{f}_{Q}\mathfrak{m}_{R,\mathsf{y}}^{[1]})^{1/2}\quad\text{ weakly in $L^{2}(\Lambda_{R}^{[1]})$},$ (5.14)
$\lim_{\mathfrak{n}\to\infty}(\mathsf{f}_{Q}\mathfrak{m}_{R,\mathsf{y}^{N_{\mathfrak{n}}}}^{N_{\mathfrak{n}},[1]})^{1/2}=(\mathsf{f}_{Q}\mathfrak{m}_{R,\mathsf{y}}^{[1]})^{1/2}\quad\text{ in $L^{2}(\Lambda_{R}^{[1]})$}.$ (5.15)
Proof.
Lemma 5.5 corresponds to Lemma 4.8 and follows from Lemma 5.2, Lemma 5.3, and Lemma 5.4 using the same argument as for the derivation of Lemma 4.8 from Lemma 4.5, Lemma 4.6, and Lemma 4.7. Hence, the proof is omitted here. ∎
We now construct the logarithmic derivative of $\mu_{R,\mathsf{y}}$.
Theorem 5.1.
Assume ((A1))((A3)). Then, for $\mu$-a.s. $\mathsf{y}$, the following hold.
i Let $R\geq Q+1$. For all $\mathscr{F}_{R}^{[1]}$-measurable $\mathsf{h}=(\mathsf{f}_{Q}\hat{\mathsf{f}})\otimes\mathsf{g}\in\mathscr{D}_{\circ\mathrm{b}}^{[1]}$ such that $\hat{\mathsf{f}}\in C_{0}^{\infty}(S_{R})$,
[…]
$\int_{\mathbb{R}^{d}\times\mathsf{S}}\mathsf{h}\mathsf{d}_{R,\mathsf{y}}d\mu_{R,\mathsf{y}}^{[1]}=-\int_{\mathbb{R}^{d}\times\mathsf{S}}\nabla_{{x}}\mathsf{h}d\mu_{R,\mathsf{y}}^{[1]}.$ (5.18)
ii For each $R\in\mathbb{N}$, $\mathsf{d}_{R,\mathsf{y}}$ is the logarithmic derivative of $\mu_{R,\mathsf{y}}$.
Proof.
Note that Lemma 5.2–Lemma 5.5 correspond to Lemma 4.5–Lemma 4.8, respectively. Hence, Theorem 5.1 follows from Lemma 5.2–Lemma 5.5 using similar arguments as for the proof of Theorem 4.2. Hence, the proof is omitted here. ∎
Remark 5.1.
In the definition of the logarithmic derivative, the space of test functions is limited to $C_{0}^{\infty}(S_{R})\otimes\mathscr{D}_{\circ\mathrm{b}}$, which is a subset of $\mathscr{D}_{\circ\mathrm{b}}^{[1]}=C_{0}^{\infty}(\mathbb{R}^{d})\otimes\mathscr{D}_{\circ\mathrm{b}}$. If we use $\mathscr{D}_{\circ\mathrm{b}}^{[1]}$ as the space of test functions, a distribution emerges from the boundary condition.
5.2. Explicit formulae of the logarithmic derivative $\mathsf{d}_{R,\mathsf{y}}$.
Let $\mathbf{I}(k)$ be as in (2.24). For $\mathbf{i}\in\mathbf{I}(k)$ and ${x}=({x}_{1},\ldots,{x}_{d})$, we set ${x}^{\mathbf{i}}={x}_{1}^{i_{1}}\cdots{x}_{d}^{i_{d}}$. Let $f^{(\mathbf{i})}=\partial^{k}f/\partial{x}^{\mathbf{i}}$, where $\partial{x}^{\mathbf{i}}=(\partial x_{1})^{i_{1}}\cdots(\partial x_{d})^{i_{d}}$.
Lemma 5.6.
Let $\ell$ be as in ((A3)). We set $\mathsf{y}=\sum_{i}\delta_{{y}^{i}}$. Then,
$\max_{\mathbf{i}\in\mathbf{I}(\ell+1)}\int_{\mathsf{S}}\sum_{{y}^{i}\notin S_{R+\epsilon}}^{\infty}\big{\lVert}(-\nabla\Psi)^{(\mathbf{i})}({x}-{y}^{i})\big{\rVert}_{C^{1}(\overline{S}_{R})}d\mu(\mathsf{y})<\infty.$ (5.19)
Proof.
According to (1.1) and $\mathbf{i}\in\mathbf{I}(\ell+1)$, (5.19) follows from
$\int_{\mathsf{S}}\sum_{{y}^{i}\notin S_{R+\epsilon}}^{\infty}\sup_{{x}\in\overline{S}_{R}}\frac{1}{\lvert{x}-{y}^{i}\rvert^{d+\ell}}d\mu(\mathsf{y})<\infty.$ (5.20)
By ((A3)), $\ell>\ell_{0}-d$. Hence, from (2.22) and (5.20), we obtain (5.19). ∎
Let $\mathbb{I}(\ell)=\sqcup_{k=0}^{\ell}\mathbf{I}(k)$ as in Theorem 2.1. Recall that, from (2.14), $\nabla_{{x}}\Psi^{N}({x},{y})=0$ for $({x},{y})\notin\overline{O_{N}\times O_{N}}$. Using Taylor’s expansion of $\nabla_{{x}}\Psi^{N_{\mathfrak{n}}}({x},{y})$ and $\nabla\Psi({x}-{y})$ in x{x} at the origin x=0{x}=0 for each |y|>R|{y}|>R and $({x},{y})\in O_{N}\times O_{N}$, we obtain, for $|{x}|\leq R$,
$-\nabla\Psi({x}-{y})$ $=\sum_{\mathbf{i}\in\mathbb{I}(\ell)}\frac{{x}^{\mathbf{i}}}{i_{1}!\cdots i_{d}!}(-\nabla\Psi)^{(\mathbf{i})}(0-{y})+\mathscr{Q}^{\ell}({x},{y}),$ (5.21)
where $\mathscr{Q}^{\ell,\mathfrak{n}}$ and $\mathscr{Q}^{\ell}$ are the residual terms of Taylor’s expansion defined by
$\mathscr{Q}^{\ell,\mathfrak{n}}({x},{y})$ $=\sum_{\mathbf{i}\in\mathbf{I}(\ell+1)}\frac{{x}^{\mathbf{i}}}{i_{1}!\cdots i_{d}!}(-\nabla_{{x}}\Psi^{N_{\mathfrak{n}}})^{(\mathbf{i})}(\xi_{{x}}^{\mathfrak{n}},{y}),$
$\mathscr{Q}^{\ell}({x},{y})$ $=\sum_{\mathbf{i}\in\mathbf{I}(\ell+1)}\frac{{x}^{\mathbf{i}}}{i_{1}!\cdots i_{d}!}(-\nabla\Psi)^{(\mathbf{i})}(\xi_{{x}}-{y}),$ (5.22)
where $\xi_{{x}}^{\mathfrak{n}}$ and $\xi_{{x}}\in S_{R}$ are such that $\lvert\xi_{{x}}^{\mathfrak{n}}\rvert,\lvert\xi_{{x}}\rvert<\lvert{x}\rvert$. Using (2.26), we can and do take $\lim_{\mathfrak{n}\to\infty}\xi_{{x}}^{\mathfrak{n}}=\xi_{{x}}$ in $C^{1}(\overline{S}_{R})$ and $\mathscr{Q}^{\ell,\mathfrak{n}}({x},{y}),\ \mathscr{Q}^{\ell}({x},{y})\in C^{1}(\overline{S}_{R})$ for each |y|>R|{y}|>R.
Let $\mathscr{R}_{R,\mathsf{y}}^{\ell,\mathfrak{n},\mathfrak{n}}$, $\mathscr{R}_{R,\mathsf{y}}^{\ell,\mathfrak{n}}$, $\mathscr{R}_{R,\mathsf{y}}^{\ell}$, and $\mathscr{R}_{R+\epsilon,\mathsf{y}}^{\ell}$ be such that, for $\mu$-a.s. $\mathsf{y}=\sum_{i}\delta_{{y}^{i}}$,
$\quad\mathscr{R}_{R,\mathsf{y}}^{\ell}({x})$ $=\sum_{{y}^{i}\not\in\overline{S}_{R}}^{\infty}\mathscr{Q}^{\ell}({x},{y}^{i}),\quad\mathscr{R}_{R+\epsilon,\mathsf{y}}^{\ell}({x})=\sum_{{y}^{i}\notin\overline{\mathsf{S}}_{R+\epsilon}}^{\infty}\mathscr{Q}^{\ell}({x},{y}^{i}).$ (5.23)
Proposition 5.1.
i For each $R\in\mathbb{N}$ and $\mu$-a.s. $\mathsf{y}=\sum_{i}\delta_{{y}^{i}}$, we obtain
$\mathscr{R}_{R,\mathsf{y}}^{\ell}=\mathscr{R}_{R,\pi_{\overline{S}_{R}^{c}}(\mathsf{y})}^{\ell},$ (5.25)
$\lim_{R\to\infty}\lVert\mathscr{R}_{R,\mathsf{y}}^{\ell}\rVert_{C^{1}(\overline{S}_{Q})}=0\quad\text{ for each $Q\in\mathbb{N}$}.$ (5.26)
ii For each $R\in\mathbb{N}$ and $\epsilon>0$,
$\int_{\mathsf{S}}\lVert\mathscr{R}_{R+\epsilon,\mathsf{y}}^{\ell}({x})\rVert_{C^{1}(\overline{S}_{R})}d\mu(\mathsf{y})<\infty.$
Proof.
Let $\mathsf{y}=\sum_{i}\delta_{{y}^{i}}$. From (5.22) and (2.27), we have, for each $\epsilon>0$,
$\int_{\mathsf{S}}\sum_{{y}^{i}\notin S_{R+\epsilon}}^{\infty}\sup_{\mathfrak{n}\in\mathbb{N}}\lVert\mathscr{Q}^{\ell,\mathfrak{n}}({x},{y}^{i})\rVert_{C^{1}(\overline{S}_{R})}d\mu(\mathsf{y})$
[…]
Hence, for $\mu$-a.s. $\mathsf{y}=\sum_{i}\delta_{{y}^{i}}$ and each $\epsilon>0$, we have
$\sum_{{y}^{i}\notin S_{R+\epsilon}}^{\infty}\sup_{\mathfrak{n}\in\mathbb{N}}\lVert\mathscr{Q}^{\ell,\mathfrak{n}}({x},{y}^{i})\rVert_{C^{1}(\overline{S}_{R})}<\infty.$ (5.27)
Because (5.27) holds for all $\epsilon>0$, we deduce that, for $\mu$-a.s. $\mathsf{y}=\sum_{i}\delta_{{y}^{i}}$,
$\sum_{{y}^{i}\not\in\overline{S}_{R}}^{\infty}\sup_{\mathfrak{n}\in\mathbb{N}}\lVert\mathscr{Q}^{\ell,\mathfrak{n}}({x},{y}^{i})\rVert_{C^{1}(\overline{S}_{R})}<\infty.$ (5.28)
From (5.28), the infinite sum $\lim_{\mathfrak{n}\to\infty}\mathscr{R}_{R,\mathsf{y}}^{\ell,\mathfrak{n},\mathfrak{n}}$ converges absolutely in $C^{1}(\overline{S}_{R})$.
From (5.22) and (5.19), we obtain, for each $\epsilon>0$,
$\int_{\mathsf{S}}\sum_{{y}^{i}\notin S_{R+\epsilon}}^{\infty}\lVert\mathscr{Q}^{\ell}({x},{y}^{i})\rVert_{C^{1}(\overline{S}_{R})}d\mu(\mathsf{y})$
[…]
$\leq\sum_{\mathbf{i}\in\mathbf{I}(\ell+1)}\frac{R^{\ell+1}}{i_{1}!\cdots i_{d}!}\int_{\mathsf{S}}\sum_{{y}^{i}\notin S_{R+\epsilon}}^{\infty}\big{\lVert}(-\nabla\Psi)^{(\mathbf{i})}({x}-{y}^{i})\big{\rVert}_{C^{1}(\overline{S}_{R})}d\mu(\mathsf{y})$
[…]
Hence, similarly as (5.28), we deduce, for $\mu$-a.s. $\mathsf{y}=\sum_{i}\delta_{{y}^{i}}$,
$\sum_{{y}^{i}\not\in\overline{S}_{R}}^{\infty}\lVert\mathscr{Q}^{\ell}({x},{y}^{i})\rVert_{C^{1}(\overline{S}_{R})}<\infty,$ (5.30)
which yields, for $\mu$-a.s. $\mathsf{y}=\sum_{i}\delta_{{y}^{i}}$,
$\lim_{R\to\infty}\sum_{{y}^{i}\not\in\overline{S}_{R}}^{\infty}\lVert\mathscr{Q}^{\ell}({x},{y}^{i})\rVert_{C^{1}(\overline{S}_{Q})}=0.$
This implies (5.26). From (5.30), the infinite sums $\mathscr{R}_{R,\mathsf{y}}^{\ell}$ and $\lim_{\mathfrak{n}\to\infty}\mathscr{R}_{R,\mathsf{y}}^{\ell,\mathfrak{n}}$ converge absolutely in $C^{1}(\overline{S}_{R})$, which implies the first equality in (5.24).
Using (5.23), (5.22), (2.29), and $\lim_{\mathfrak{n}\to\infty}\xi_{{x}}^{\mathfrak{n}}=\xi_{{x}}$, we have
$\lim_{\mathfrak{n}\to\infty}\lVert\mathscr{R}_{R,\mathsf{y}}^{\ell,\mathfrak{n},\mathfrak{n}}({x})-\mathscr{R}_{R,\mathsf{y}}^{\ell,\mathfrak{n}}({x})\rVert_{C^{1}(\overline{S}_{R})}$
[…]
$\leq\lim_{\mathfrak{n}\to\infty}\sum_{\mathbf{i}\in\mathbf{I}(\ell+1)}\frac{{x}^{\mathbf{i}}}{i_{1}!\cdots i_{d}!}\sum_{{y}^{i}\not\in\overline{S}_{R}}^{\infty}\lVert(-\nabla\Psi^{N_{\mathfrak{n}}})^{(\mathbf{i})}(\xi_{{x}}^{\mathfrak{n}},{y}^{i})-(-\nabla\Psi)^{(\mathbf{i})}(\xi_{{x}}-{y}^{i})\rVert_{C^{1}(\overline{S}_{R})}$
[…]
Here, we used (5.22) for the third line. (5.31) yields the second equality in (5.24).
From (5.23), $\mathscr{R}_{R,\mathsf{y}}^{\ell,\mathfrak{n}}$ is independent of $\pi_{\overline{S}_{R}}(\mathsf{y})$. Hence, (5.24) yields (5.25). Thus, we have obtained i.
From (5.23) and (5.29), we have
$\int_{\mathsf{S}}\lVert\mathscr{R}_{R+\epsilon,\mathsf{y}}^{\ell}({x})\rVert_{C^{1}(\overline{S}_{R})}d\mu(\mathsf{y})\leq\int_{\mathsf{S}}\sum_{{y}^{i}\notin S_{R+\epsilon}}^{\infty}\lVert\mathscr{Q}^{\ell}({x},{y}^{i})\rVert_{C^{1}(\overline{S}_{R})}d\mu(\mathsf{y})<\infty.$
This implies ii. ∎
For $\mathbf{i}=(i_{1},\ldots,i_{d})\in\sqcup_{k=0}^{\infty}\mathbf{I}(k)$ and $\mathsf{y}=\sum_{i}\delta_{{y}^{i}}$, we set
Let $\mathbb{I}(\ell)=\sqcup_{k=0}^{\ell}\mathbf{I}(k)$ as in Theorem 2.1. From (5.21)–(5.23) and (5.32),
Proposition 5.2.
For each $\mathbf{i}\in\mathbb{I}(\ell)$, $\mu$-a.s. $\mathsf{y}=\sum_{i}\delta_{{y}^{i}}$ such that $\mathsf{y}(\overline{S}_{R})\geq 1$, and each $R\in\mathbb{N}$, we have a finite limit $\eta_{R,\mathsf{y}}^{\mathbf{i}}$ such that
$\lim_{R\to\infty}\eta_{R,\mathsf{y}}^{\mathbf{i}}=0.$ (5.37)
Proof.
We first prove the existence of a finite $\eta_{R,\mathsf{y}}^{\mathbf{i}}$. From (5.33) and (5.5),

### v4

@@ 10.1. Tail decomposition of Dirichlet forms and weak solution
$\underline{\mathbf{X}}_{\mathsf{a}}^{[m]}$ $=\lim_{R\to\infty}\underline{\mathbf{X}}_{\mathsf{a},R}^{[m]}\quad\text{ in law in $C([0,\infty);(\mathbb{R}^{d})^{m}\times\mathsf{S})$},$
$\underline{\mathbf{X}}_{\mathsf{a}}$ $=\lim_{R\to\infty}\underline{\mathbf{X}}_{\mathsf{a},R}\quad\text{ in law in $W(\mathbb{R}^{d})^{\mathbb{N}}$}.$
In particular, each tagged particle $\underline{X}_{\mathsf{a}}^{i}$ of $\underline{\mathbf{X}}_{\mathsf{a}}$ does not explode.
Proof.
The proof is the same as that of Theorem 8.1 and is omitted. ∎
Proposition 10.2.
For $\mu$-a.s. $\mathsf{a}$, the following hold.
i There exists a $\mu_{\mathsf{a}}^{[m]}$-symmetric, conservative diffusion $\overline{\mathbf{X}}_{\mathsf{a}}^{[m]}$ properly associated with $(\mathscr{E}_{\mathsf{a}}^{[m]},\mathscr{D}_{\mathsf{a},\mathrm{upr}}^{[m]})$ on $L^{2}(\mu_{\mathsf{a}}^{[m]})$ for each $m\in\{0\}\cup\mathbb{N}$.
where $r_{t}$ is the time reversal operator defined before Lemma 8.1.
Proof.
The proof is similar to that of Theorem 9.1. Hence, it is omitted. ∎
Let $\overline{\mathbf{X}}_{\mathsf{a}}$ be the fully labeled process constructed in the same fashion as $\overline{\mathbf{X}}$ in Theorem 9.2. Then $\overline{\mathbf{X}}_{\mathsf{a}}^{[m]}$ in Theorem 10.2 is the $m$-labeled process of $\overline{\mathbf{X}}_{\mathsf{a}}$.
Proposition 10.3.
[…]
Proof.
The proof of Proposition 10.3 is identical to the combination of those for Theorems 8.2 and 9.2. Hence, it is omitted. ∎
10.2. Unique strong solution: Proof of Theorem 2.3
@@
Let $\mathsf{K}[a_{\mathsf{q}}]$ be the compact set defined in (13.4). We can easily find a sequence $\mathbf{a}=\{a_{\mathsf{q}}\}_{\mathsf{q}\in\mathbb{N}}$ as in (13.3) satisfying (13.5) and
$\sum_{\mathsf{q}=1}^{\infty}\mathsf{q}^{2}\mu_{\mathsf{a}}(\mathsf{K}[a_{\mathsf{q}}]^{c})<\infty.$ (10.2)
@@
We take a $\mu_{\mathsf{a}}^{[1]}$-version of $\mathsf{d}^{\mu_{\mathsf{a}}}$ as in Theorem 2.1 and Proposition 5.3. Here $\mathsf{y}$ in Proposition 5.3 is replaced by $\mathsf{a}$. Then, for $\mu_{\mathsf{a}}$-a.s. $\mathsf{s}$, $\mathsf{d}^{\mu_{\mathsf{a}}}({x},\mathsf{s})$ is locally Lipschitz in x{x} on $\mathbb{R}^{d}_{\neq}(\mathsf{s})$. Let $\mathsf{S}_{\neq}^{[1]}[\mathsf{a}]$ be as in (5.36). Let
$\mathsf{S}_{\mathrm{sde}}=\mathfrak{u}^{[1,0]}(\mathsf{S}_{\neq}^{[1]}[\mathsf{a}]),$ (10.3)
@@
In what follows, we denote both $\underline{\mathbf{X}}_{\mathsf{a}}$ and $\overline{\mathbf{X}}_{\mathsf{a}}$ by the same symbol $\mathbf{X}$, and denote by $\mathsf{X}$ the unlabeled dynamics associated with $\mathbf{X}$.
Let $\mathsf{S}_{\mathrm{s}}=\{\mathsf{s}\in\mathsf{S};\mathsf{s}(\{{x}\})\in\{0,1\}\text{ for all }{x}\in\mathbb{R}^{d}\}$.
According to (13.9)–(13.11), condition ((B1)) iii follows from
$P\bigl(\lim_{\mathsf{q}\to\infty}\kappa_{\mathsf{q}}=\infty\bigr)=1.$ (10.5)
We have already proved (10.4) in Proposition 6.1. Thus, it only remains to prove (10.5). To verify (10.5), we use Lemma 13.9.